OpenStudy (anonymous):

I need help (in steps) to solve this: 4x(x-25)=25(1-4x)

OpenStudy (anonymous):

distribute 4x(x - 25) and then distribute 25(1-4x)

OpenStudy (anonymous):

(4x * 1x) + (4x * -25) = (25 * 1) + (25 * 4x) 4x + -100x = 25 + 100x

OpenStudy (anonymous):

So... 4x^2-100=25-100x

OpenStudy (anonymous):

yes chels :D

OpenStudy (anonymous):

now we combine like terms :P

OpenStudy (anonymous):